Find the general solution of the equation y''− 4y'+ 13y = 0.


1
Expert's answer
2021-09-06T16:32:21-0400

Let us find the general solution of the differential equation "y''\u2212 4y'+ 13y = 0". The characteristic equation "k^2-4k+13=0" is equivalent to "(k-2)^2=-9," and hence has the roots "k_1=2+3i" and "k_2=2-3i." Therefore, the general solutions is of the form:

"y=e^{2x}(C_1\cos (3x)+C_2\sin(3x))."